Find the measure of an angle whose measure is 18 degrees less than one-half of the measure of its complement.
Answer by stanbon(75887)

Find the measure of an angle whose measure is 18 degrees less than one-half of the measure of its complement.
-------------
angle: x
complement: 90-x
---
x = (1/2)(90-x)-18
------------
x = 45-(x/2)-18
---
(3/2)x = 27
x = (2/3)(27)
x = 18
==============
Cheers,
Stan H.
============